### v1

@@ 3.1.2 Visual Feature Extraction
The feature extraction process involves the utilization of the ResNet50 architecture, selectively extracting features up to the ‘$conv4\_block6\_3\_conv$’ layer. The intention behind this decision is to capture complex and abstract representations that reside in deeper layers of the network. This is additionally performed to ensure that the shapes align with those of the text encoder, facilitating the interaction of features through the IDEA attention mechanism. The residual connections in ResNet50 facilitate effective gradient flow during training, addressing the vanishing gradient problem. The resulting feature map ($F_{\text{emb}}$) encapsulates hierarchical information, ranging from fundamental attributes such as edges and textures to more complex features such as object components and scenes. A bespoke convolutional layer, denoted as ‘$custom\_conv$’, is seamlessly integrated into the architecture. This custom convolutional layer introduces 1024 filters and employs a (4, 4) kernel, inducing a transformative effect on the output shape, resulting in dimensions (11, 11, 1024). Subsequently, a crucial reshaping operation is applied to the feature map, yielding a tensor with dimensions (1, 121, 1024). This reshaping is strategically aligned with the subsequent processing requirements, ensuring optimal compatibility with the changing tensor structure. Simultaneously, a padding operation is introduced to the tensor to retain spatial information and maintain dimensional consistency. This step results in a tensor shape of (1, 128, 1024), crucial for the seamless progression of information through the network architecture. To safeguard the knowledge encoded in the ResNet50 pre-trained weights and fine-tune the model for the specialized image processing task, a discerning decision is made to freeze all original layers. This strategic freezing confines the training scope exclusively to the parameters of the ‘$custom\_conv$’ layer. Equation (2) details the calculation of each element in the final visual feature matrix $H_{v}$, where the variable $i$ represents the index of the features in the final visual feature matrix ${H_{v}}$. $H_{\text{img}}$ and $W_{\text{img}}$ represent the height and width of the feature map, respectively, and $k$ denotes a specific channel within $F_{\text{emb}}$.
@@
3.1.3 Inverted Dual Embedded Attention (IDEA)
In multimodal data analysis, we often face the challenge of efficiently integrating information from various modalities, such as text and image, to achieve a holistic comprehension of the underlying content. Traditional attention mechanisms focus on aggregating information from multiple sources in which the modalities are aligned. However, they don’t focus on the contrary information present in the modalities. We propose an approach that recognizes both harmonious and contrary information present in the different modalities. The contrary information present between the modalities should be identified for several reasons. Contrary pairs highlight differences and inconsistencies between modalities. They contribute to enhanced model interpretability and provide insights into why a model makes specific decisions. They are crucial for tackling problems with noisy or misaligned multimodal data, ensuring a stronger and more dependable analysis. In this context, we introduce an attention mechanism called Inverted Dual Embedded Attention (IDEA). It combines harmonious and contrary attention vectors to reveal the complex interactions between different types of data. This unique strategy not only enhances the interpretability of the model but also provides a deeper understanding of multimodal information. By considering both harmonious and contrary attention, IDEA adapts to varying information needs, facilitating richer insights, improved representational learning, and a versatile approach to multimodal data analysis. IDEA contains four modules: (a) Shared Embedding Module; (b) Harmonious Attention Module; (c) Contrary Attention Module; and (d) Attented Fusion Module.
@@ 3.1.3.1.1 Textual Modality Transformation:
In the initial stage, textual features $H_{t}\in\mathbb{R}^{d\times dt}$ undergo a non-linear projection into a shared embedding space. This involves a linear transformation using trainable parameters $W_{t}\in\mathbb{R}^{d_{t}\times d_{se}}$ and $b_{t}\in\mathbb{R}^{1\times d_{se}}$, resulting in $H_{t}W_{t}+b_{t}$ as the linearly projected textual features within the shared embedding space. The linearly projected textual features are subjected to a non-linear transformation denoted as $H_{{Text}}\in\mathbb{R}^{d\times d_{se}}$ that includes batch normalization [38, 39] and a hyperbolic tangent activation. Batch normalization helps in stabilizing and accelerating the training of deep neural networks by normalizing the inputs to each layer, reducing internal covariate shifts, and accelerating the training of deep neural networks. It normalizes the input of a layer by adjusting and scaling the activations. This can be shown in Eq. 3:
@@ 3.1.3.1.2 Visual Modality Transformation:
The linearly projected visual features are subjected to a non-linear transformation denoted as $H_{{Vis}}\in\mathbb{R}^{d\times d_{se}}$ that includes batch normalization [38, 39] and a hyperbolic tangent activation as expressed by the Eq. 4:
@@
where $\gamma$, $\beta$, $\mu$ and $\sigma$ are already defined in Eq. 3 The dimensions of various parameters in our case are $d=128$, $d_{se}=1024$, $d_{t}=768$, and $d_{v}=1024$.
3.1.3.1.3 Similarity Matrix Computation:
To compute the similarity matrix $S_{{sim}}[i,j]\in\mathbb{R}^{d\times d}$, we first compute $\hat{S}_{\text{sim}}[i,k]\in\mathbb{R}^{d\times d_{se}}$ by applying the element-wise tangent hyperbolic (tanh) function to the sum of the projected textual feature vectors $H_{\text{Text}}[i,m]\in\mathbb{R}^{d\times d_{se}}$ and visual feature vectors $H_{\text{Vis}}[i,m]\in\mathbb{R}^{d\times d_{se}}$ as shown in the Eq. 5.
$\hat{S}_{\text{sim}}[i,k]=\tanh\left(\sum_{m=1}^{d_{\text{se}}}\left(H_{\text{Vis}}[i,m]+H_{\text{Text}}[i,m]\right)\right)$ (5)
Here, $i$ is the index that ranges over the $d$ dimensions of the feature vectors, and $k$ is the index for the $d_{se}$ dimensions of the shared space. The index $m$ is used for the summation, which ranges from 1 to $d_{se}$. Each element $\hat{S}_{\text{sim}}[i,k]$ of the resulting matrix represents the $\tanh$ of the sum of the $m$-th elements of the $i$-th row of $H_{\text{Vis}}$ and $H_{\text{Text}}$.
Subsequently, a linear transformation is applied using the trainable weight matrix $W_{\text{sim}}[k,j]\in\mathbb{R}^{d_{{se}}\times d}$ to learn the importance of different elements in the similarity computation. The final similarity matrix $S_{\text{sim}}$ can be expressed as shown in the Eq. 6.
$S_{\text{sim}}[i,j]=\sum_{k=1}^{d_{\text{se}}}\hat{S}_{\text{sim}}[i,k]\cdot W_{\text{sim}}[k,j]$ (6)
In the similarity matrix $S_{\text{sim}}[i,j]\in\mathbb{R}^{d\times d}$, every element in row $i$ captures the degree of similarity between the $i$-th visual feature and the entire set of textual features. Correspondingly, each element in column $j$ quantifies the similarity between the $j$-th textual feature and the entire set of visual features.
Recent research studies [40, 41] have predominantly focused on using attention matrices to evaluate co-attention weights by comparing similarities among different modalities. In contrast, our approach, “Inverted Dual Embedded Attention (IDEA)”, does things differently. It calculates attention in a way that highlights differences and similarities, helping us understand how individual aspects of one type of information relate to the whole set of aspects of another type. It captures the hidden and complex information via two attention mechanisms, the Harmonious Attention Mechanism (HAM) and the Contrary Attention Mechanism (CAM), as shown in Fig. 2.
Figure 2: Inverted Dual Embedded Attention (IDEA)
@@ 3.1.3.2 Harmonious Attention Mechanism (HAM):
Harmonious attention mechanism emphasizes the alignment and correlation between features or information from different modalities. It is used to identify and highlight shared or similar aspects of the data. In a multimodal sentiment analysis task, harmonious attention is defined as the part that looks at words in a text description and elements in an accompanying image that express the same emotional sentiment. This aligns textual and visual cues associated with sentiment effectively. The harmonious attention vector facilitates the identification of highly correlated features between the two modalities. So, in a harmonious attention mechanism, we obtain the visual-guided textual features and vice versa. The visual-enriched textual features are obtained by applying the softmax function to the similarity matrix. This process yields attention weights, which are subsequently multiplied by the textual features. We introduce a temperature parameter $T$ in the softmax function to control the softness or sharpness of the probability distribution[42]. A higher temperature encourages a softer, more uniform probability distribution, while a lower temperature value results in a sharper, more peaked probability distribution [43, 44]. For HAM, a higher temperature might lead to more evenly spread attention over the elements of the similarity matrix, whereas a lower temperature might result in more focused attention on a few specific elements. After rigorous experimentation, the value of T for HAM is 1.65, and for CAM, it is 0.75. We obtain the attention weights or scores as expressed in Eq. 7:
@@
$Att_{\text{weights}}$ is a matrix where each element $(Att_{\text{weights}})_{i,j}$ represents the attention weight for the corresponding element in the similarity matrix $S_{\text{sim}}$ along axis $i$. The visual-enriched textual features $T_{\text{HAM}}\in\mathbb{R}^{d\times dt}$ can be computed as shown in the Eq. 8
@@
Similarly, the textual-enriched visual features $V_{\text{HAM}}\in\mathbb{R}^{d\times d_{v}}$ are computed by employing the softmax function on the similarity matrix to produce attention weights, which are then multiplied by the visual features. Since the attention weights $(Att_{\text{weights}})_{i,j}$ have already been computed in Eq. 7, we will use the same attention weights for the computation of $V_{\text{HAM}}$ as shown in Eq. 9.
@@ 3.1.3.3 Contrary Attention Mechanism (CAM):
Contrary attention highlights differences and disparities between features or information from different modalities. It is used to identify and accentuate areas where modalities diverge or provide contrary information. In a multimodal content moderation task, contrary attention might identify text and image content that contradict each other, helping the model detect potentially divergent content where the text and visual elements are in conflict. However, in most prior studies, contrary information, which can also be valuable for representation, is often overlooked. As shown in Fig. 2, we compute CAM using the similarity matrix and the text and image modalities after multiplying them by -1 (designated as Invert). Multiplying embeddings with -1 will reverse the direction of the embedding in the vector space. It does not alter the magnitude of the features. This can be useful for tasks such as learning to distinguish between real and fake data, generating adversarial examples, and finding various different hidden patterns. The inverted visual-enriched textual features $T_{\text{CAM}}\in\mathbb{R}^{d\times dt}$ are computed as shown in Eq. 10.
@@
Similarly, the inverted textual-enriched visual features $V_{\text{CAM}}\in\mathbb{R}^{d\times dv}$ are computed as shown in Eq. 11:
@@ 3.1.3.4 Fused Attended Module:
To combine both harmonious and contrary attention vectors, we utilize dense layers akin to the structure found in multi-head attention [45], as illustrated below: The textually-informed visual features are computed as shown in Eq. 12:
@@
where $T_{\text{fuse}}\in\mathbb{R}^{d\times dt}$, $W_{\text{text}}\in\mathbb{R}^{2dt\times dt}$ and $b_{\text{text}}\in\mathbb{R}^{1\times dt}$ are trainable parameters. After obtaining $T_{\text{fuse}}$ and $V_{\text{fuse}}$, We utilize dense layers to capture and understand the relationship that exists between the visual features and the enriched fused features. This process enables us to obtain what we refer to as the visual-fusion feature ($V_{\text{fusion}}$), as shown in Eq. 14.
$V_{fusion}=tanh((T_{fuse}\oplus H_{v})\widetilde{W_{v}}+\widetilde{b_{v}})$ (14)
@@
Similarly, to obtain the textual-fusion feature ($T_{fusion}$), we employ dense layers to capture the correlation between the textual features and the enhanced fused features, as shown in Eq. 15.
$T_{fusion}=tanh((V_{fuse}\oplus H_{t})\widetilde{W_{t}}+\widetilde{b_{t}})$ (15)
@@
After obtaining $V_{\text{fusion}}$ and $T_{\text{fusion}}$, the next step involves extracting the most significant representation from the combined textual and visual features. We apply global average pooling to the first dimension of $T_{\text{fusion}}\in\mathbb{R}^{d\times dt}$ and $V_{\text{fusion}}\in\mathbb{R}^{d\times dv}$, the dimensions reduce to $T_{\text{fusion}}\in\mathbb{R}^{1\times dt}$ and $V_{\text{fusion}}\in\mathbb{R}^{1\times dv}$ respectively. It is used to reduce dimensionality, prevent overfitting, and provide a summarized representation of the feature maps. We compute the attention scores on the pooled features using a self-attention mechanism [46].
@@
In Eq. 16, $W_{Q}\in\mathbb{R}^{dt\times dz}$ and $W_{k}\in\mathbb{R}^{dt\times dz}$ are parameter matrices. The term $\sqrt{\text{dz}}$ represents the scaling factor, where dz is the dimension of the query and key spaces.
The attention weight ($\alpha_{i}$) is obtained by applying the softmax function to the attention scores as shown in  Eq. 17:
@@
The final textual fused feature vector representation $T_{\text{final}}\in\mathbb{R}^{1\times dt}$ is a weighted sum of the values based on attention weights as shown in Eq. 18:
@@
Similarly, for $V_{\text{fusion}}$, repeat Eqs. 16, 17 and 18, by replacing $T_{\text{fusion}}$ with $V_{\text{fusion}}$, we can get the final representation of visual fused feature vector as $V_{\text{final}}\in\mathbb{R}^{1\times dv}$.
After obtaining the final textual-fused feature vector $T_{\text{final}}$ through the attention mechanism, the last step involves concatenating it with the final visual-fused feature vector $V_{\text{final}}$. This results in a single fused attended vector (FAV)$\in\mathbb{R}^{1\times(dt+dv)}$ that captures the most significant representations from both textual and visual features. The concatenation operation is shown in Eq. 19.

### v2

@@ 4.1.2 Visual Feature Extraction
The feature extraction process involves the utilization of the ResNet50 architecture, selectively extracting features up to the ‘conv4_block6_3_convconv4\_block6\_3\_conv’ layer. The intention behind this decision is to capture complex and abstract representations that reside in deeper layers of the network. This is additionally performed to ensure that the shapes align with those of the text encoder, facilitating the interaction of features through the IDEA attention mechanism. The residual connections in ResNet50 facilitate effective gradient flow during training, addressing the vanishing gradient problem. The resulting feature map ($F_{\text{emb}}$) encapsulates hierarchical information, ranging from fundamental attributes such as edges and textures to more complex features such as object components and scenes. A bespoke convolutional layer, denoted as ‘custom_convcustom\_conv’, is seamlessly integrated into the architecture. This custom convolutional layer introduces 1024 filters and employs a (4, 4) kernel, inducing a transformative effect on the output shape, resulting in dimensions (11, 11, 1024). Subsequently, a crucial reshaping operation is applied to the feature map, yielding a tensor with dimensions (1, 121, 1024). This reshaping is strategically aligned with the subsequent processing requirements, ensuring optimal compatibility with the changing tensor structure. Simultaneously, a padding operation is introduced to the tensor to retain spatial information and maintain dimensional consistency. This step results in a tensor shape of (1, 128, 1024), crucial for the seamless progression of information through the network architecture. To safeguard the knowledge encoded in the ResNet50 pre-trained weights and fine-tune the model for the specialized image processing task, a discerning decision is made to freeze all original layers. This strategic freezing confines the training scope exclusively to the parameters of the ‘custom_convcustom\_conv’ layer. Equation (4) details the calculation of each element in the final visual feature matrix $H_{v}$, where the variable $i$ represents the index of the features in the final visual feature matrix $H_{v}$}. $H_{\text{img}}$ and $W_{\text{img}}$ represent the height and width of the feature map, respectively, and $k$ denotes a specific channel within $F_{\text{emb}}$.
@@
4.1.3 Inverted Dual Embedded Attention
In multimodal data analysis, we often face the challenge of efficiently integrating information from various modalities, such as text and image, to achieve a holistic comprehension of the underlying content. Traditional attention mechanisms focus on aggregating information from multiple sources in which the modalities are aligned. However, they don’t focus on the contrary information present in the modalities. We propose an approach that recognizes both harmonious and contrary information present in the different modalities. The contrary information present between the modalities should be identified for several reasons. Contrary pairs highlight differences and inconsistencies between modalities. They contribute to enhanced model interpretability and provide insights into why a model makes specific decisions. They are crucial for tackling problems with noisy or misaligned multimodal data, ensuring a stronger and more dependable analysis. In this context, we introduce an attention mechanism called Inverted Dual Embedded Attention (IDEA). It combines harmonious and contrary attention vectors to reveal the complex interactions between different types of data. This unique strategy not only enhances the interpretability of the model but also provides a deeper understanding of multimodal information. By considering both harmonious and contrary attention, IDEA adapts to varying information needs, facilitating richer insights, improved representational learning, and a versatile approach to multimodal data analysis. IDEA contains four modules: (a) Shared Embedding Module; (b) Harmonious Attention Module; (c) Contrary Attention Module; and (d) Attended Fusion Module.
@@ 4.1.3.1.1 Textual Modality Transformation:
In the initial stage, textual features $H_{t}\in\mathbb{R}^{d\times dt}$ undergo a non-linear projection into a shared embedding space. This involves a linear transformation using trainable parameters $W_{t}\in\mathbb{R}^{d_{t}\times d_{se}}$ and $b_{t}\in\mathbb{R}^{1\times d_{se}}$, resulting in $H_{t}W_{t}+b_{t}$ as the linearly projected textual features within the shared embedding space. The linearly projected textual features are subjected to a non-linear transformation denoted as $H_{{Text}}\in\mathbb{R}^{d\times d_{se}}$ that includes batch normalization (Luo et al., 2019; Kohler et al., 2018) and a hyperbolic tangent activation. Batch normalization helps in stabilizing and accelerating the training of deep neural networks by normalizing the inputs to each layer, reducing internal covariate shifts, and accelerating the training of deep neural networks. It normalizes the input of a layer by adjusting and scaling the activations. This can be shown in Eq.˜5:
@@ 4.1.3.1.2 Visual Modality Transformation:
The linearly projected visual features are subjected to a non-linear transformation denoted as $H_{{Vis}}\in\mathbb{R}^{d\times d_{se}}$ that includes batch normalization (Luo et al., 2019; Kohler et al., 2018) and a hyperbolic tangent activation as expressed by the Eq.˜6:
@@
where $\gamma$, $\beta$, $\mu$ and $\sigma$ are already defined in Eq.˜5 The dimensions of various parameters in our case are $d=128$, $d_{se}=1024$, $d_{t}=768$, and $d_{v}=1024$.
4.1.3.1.3 Attention Matrix Computation:
The attention matrix $S_{{sim}}[i,j]\in\mathbb{R}^{d\times d}$ is a crucial intermediary representation that quantifies the degree of interaction between elements in two sequences, such as text and image embeddings. The computation of the attention score matrix involves taking the dot product of the text embeddings matrix and the transpose of the image embeddings matrix, followed by scaling the result by the square root of the dimensionality of the shared embeddings ($d_{\text{se}}$) as shown in the Eq.˜7.
$S_{{sim}}[i,j]=\frac{{H_{\text{Text}}}\cdot{H_{\text{Vis}}}^{T}}{\sqrt{d_{\text{se}}}}$ (7)
The dot product operation captures the pairwise interactions between corresponding elements in the text and image sequences, while the scaling factor helps stabilize the softmax function during subsequent steps. The resulting attention matrix $S_{\text{sim}}[i,j]\in\mathbb{R}^{d\times d}$, every element in row $i$ captures the degree of similarity between the $i$-th visual feature and the entire set of textual features. Correspondingly, each element in column $j$ quantifies the similarity between the $j$-th textual feature and the entire set of visual features. This matrix serves as a foundation for generating attention weights and ultimately producing an attention-weighted sum of the original embeddings, facilitating the model’s focus on relevant information in both sequences. Recent research studies (Lu et al., 2016; Jiang et al., 2020) have predominantly focused on using attention matrices to evaluate co-attention weights by comparing similarities among different modalities. In contrast, our approach, “Inverted Dual Embedded Attention (IDEA)”, does things differently. It calculates attention in a way that highlights differences and similarities, helping us understand how individual aspects of one type of information relate to the whole set of aspects of another type. It captures the hidden and complex information via two attention mechanisms, the Harmonious Attention Mechanism (HAM) and the Contrary Attention Mechanism (CAM), as shown in Fig. 2.
Figure 2: Inverted Dual Embedded Attention
@@ 4.1.3.2 Harmonious Attention Mechanism (HAM):
Harmonious attention mechanism emphasizes the alignment and correlation between features or information from different modalities. It is used to identify and highlight shared or similar aspects of the data. In a multimodal sentiment analysis task, harmonious attention is defined as the part that looks at words in a text description and elements in an accompanying image that expresses the same emotional sentiment. This aligns textual and visual cues associated with sentiment effectively. The harmonious attention vector facilitates the identification of highly correlated features between the two modalities. So, in a harmonious attention mechanism, we obtain the visual-guided textual features and vice versa. The visual-enriched textual features are obtained by applying the softmax function to the attention matrix. This process yields attention weights, which are subsequently multiplied by the textual features. We introduce a temperature parameter $T$ in the softmax function to control the softness or sharpness of the probability distribution(Mozafari et al., 2019). A higher temperature encourages a softer, more uniform probability distribution, while a lower temperature value results in a sharper, more peaked probability distribution (Hinton et al., 2015; Nguyen et al., 2020). For HAM, a higher temperature might lead to more evenly spread attention over the elements of the attention matrix, whereas a lower temperature might result in more focused attention on a few specific elements. After rigorous experimentation, the value of T for HAM is 1.65, and for CAM, it is 0.75. We obtain the attention weights or scores as expressed in Eq.˜8:
@@
$Att_{\text{weights}}$ is a matrix where each element $(Att_{\text{weights}})_{i,j}$ represents the attention weight for the corresponding element in the attention matrix $S_{\text{sim}}$ along axis $i$. The visual-enriched textual features $T_{\text{HAM}}\in\mathbb{R}^{d\times dt}$ can be computed as shown in the Eq.˜9
@@
Similarly, the textual-enriched visual features $V_{\text{HAM}}\in\mathbb{R}^{d\times d_{v}}$ are computed by employing the softmax function on the attention matrix to produce attention weights, which are then multiplied by the visual features. Since the attention weights $(Att_{\text{weights}})_{i,j}$ have already been computed in Eq.˜8, we will use the same attention weights for the computation of $V_{\text{HAM}}$ as shown in Eq.˜10.
@@ 4.1.3.3 Contrary Attention Mechanism (CAM):
Contrary attention highlights differences and disparities between features or information from different modalities. It is used to identify and accentuate areas where modalities diverge or provide contrary information. In a multimodal content moderation task, contrary attention might identify text and image content that contradict each other, helping the model detect potentially divergent content where the text and visual elements are in conflict. However, in most prior studies, contrary information, which can also be valuable for representation, is often overlooked. As shown in Fig. 2, we compute CAM using the attention matrix and the text and image modalities after multiplying them by -1 (designated as Invert). Multiplying embeddings with -1 will reverse the direction of the embedding in the vector space (hin Cheung and man Lam, 2022). It does not alter the magnitude of the features. This can be useful for tasks such as learning to distinguish between real and fake data, generating adversarial examples, and finding various different hidden patterns. The inverted visual-enriched textual features $T_{\text{CAM}}\in\mathbb{R}^{d\times dt}$ are computed as shown in Eq.˜11.
@@
Similarly, the inverted textual-enriched visual features $V_{\text{CAM}}\in\mathbb{R}^{d\times dv}$ are computed as shown in Eq.˜12:
@@ 4.1.3.4 Fused Attended Module:
To combine both harmonious and contrary attention vectors, we utilize dense layers akin to the structure found in multi-head attention (Hao et al., 2017), as illustrated below: The textually-informed visual features are computed as shown in Eq.˜13:
@@
where $T_{\text{fuse}}\in\mathbb{R}^{d\times dt}$, $W_{\text{text}}\in\mathbb{R}^{2dt\times dt}$ and $b_{\text{text}}\in\mathbb{R}^{1\times dt}$ are trainable parameters. After obtaining $T_{\text{fuse}}$ and $V_{\text{fuse}}$, We utilize dense layers to capture and understand the relationship that exists between the visual features and the enriched fused features. This process enables us to obtain what we refer to as the visual-fusion feature ($V_{\text{fusion}}$), as shown in Eq.˜15.
[…]
@@
Similarly, to obtain the textual-fusion feature ($T_{fusion}$), we employ dense layers to capture the correlation between the textual features and the enhanced fused features, as shown in Eq.˜16.
[…]
@@
After obtaining $V_{\text{fusion}}$ and $T_{\text{fusion}}$, the next step involves extracting the most significant representation from the combined textual and visual features. We apply global average pooling to the first dimension of $T_{\text{fusion}}\in\mathbb{R}^{d\times dt}$ and $V_{\text{fusion}}\in\mathbb{R}^{d\times dv}$, the dimensions reduce to $T_{\text{fusion}}\in\mathbb{R}^{1\times dt}$ and $V_{\text{fusion}}\in\mathbb{R}^{1\times dv}$ respectively. It is used to reduce dimensionality, prevent overfitting, and provide a summarized representation of the feature maps. We compute the attention scores on the pooled features using a self-attention mechanism (Shaw et al., 2018).
@@
In Eq.˜17, $W_{Q}\in\mathbb{R}^{dt\times dz}$ and $W_{k}\in\mathbb{R}^{dt\times dz}$ are parameter matrices. The term $\sqrt{\text{dz}}$ represents the scaling factor, where dz is the dimension of the query and key spaces.
The attention weight ($\alpha_{i}$) is obtained by applying the softmax function to the attention scores as shown in Eq.˜18:
@@
The final textual fused feature vector representation $T_{\text{final}}\in\mathbb{R}^{1\times dt}$ is a weighted sum of the values based on attention weights as shown in Eq.˜19:
@@
Similarly, for $V_{\text{fusion}}$, repeat Eqs.˜17, 18 and 19, by replacing $T_{\text{fusion}}$ with $V_{\text{fusion}}$, we can get the final representation of visual fused feature vector as $V_{\text{final}}\in\mathbb{R}^{1\times dv}$.
[…]